Help solve -(1+7x)-6(7-x)=36

3(w-2)
--------=w-2
4

Found 2 solutions by Alan3354, Boreal:
Answer by Alan3354(69443)

Not clear.
Is that 2 problems, or 1?

Answer by Boreal(15235)

-(1+7x)-6(7-x)=36
distribute
-1-7x-42+6x=36
-43-x=36
-x=79
x=-79
-(1-553)-6(86)=36
552-516=36
================
3(w-2)/4=w-2
multiply both sides by 4 and distribute
3w-6=4w-8
add 8 and -3w to both sides
w=2
3(0)/4=0